# Gender Dynamics and Homophily in a Social Network of LLM Agents

### v1

@@ 2.2.3 Social influence
Homophily analyses examine whether agents who perform similarly are more likely to be connected, but they do not establish whether agents become more similar to those they follow over time. To study the extent of such social influence on gender performance, we estimate a panel regression model for weekly changes in gender-scores.
For each agent $i$ and week $t$, we denote by $X_{it}$ the standardised gender-score constructed from the weekly content. We then model the week-to-week change
@@
where $\alpha_{i}$ are agent fixed effects that absorb time-invariant differences in average gender performance, $\tau_{t}$ are week fixed effects that capture shocks common to all agents, and $\varepsilon_{it}$ is an error term. The coefficient $\varphi$ measures self-correction, that is, the tendency for agents to revert toward their own typical level of gendered language, while $\gamma_{\mathrm{inf}}$ captures social influence, that is, the extent to which agents shift their gender-score in the direction of their followees.
Because agents who post in similar ways are more likely to connect in the first place, the peer mean $\bar{X}_{N(i),t-1}$ may be endogenous. We therefore estimate both an ordinary least squares (OLS) version of the model and an instrumental variables (IV) version using two-stage least squares. Full details are described in ESM, Section D.
@@ 3.1 Structural features of the followership network
We begin by describing the followership network of Chirpers (Figure 2). For detailed definitions and the complete results, see ESM, Sections E and F. Social networks are typically sparse, meaning that only a small fraction of all possible ties are realised. Consistent with this, the Chirper network remains low in density throughout most of the year, although density rises from the middle of the period onward as agents create additional follow ties. Reciprocity also increases substantially over time, so that a growing fraction of follow relations are mutual, suggesting the gradual formation of more stable bidirectional relations among LLM agents.
Figure 2: Followership network for weeks 4 and 35. The visualisations are made using ForceAtlas2 layout on Gephi $0.1$, with nooverlap option.
The degree distribution of the network is highly heterogeneous, with a minority of agents attracting many more followers than most others. Early to middle in the observation window, the in-degree and out-degree distributions on logarithmic axes are close to the pattern expected under power law behaviour. Later in the year, and especially by week 52, the distributions bend away from a pure power law, with a broad peak at intermediate degrees and a sharp drop in the upper tail. This pattern is consistent with previous work showing that growing networks often move from scale-free degree distributions toward lognormal or mixed forms as constraints on attention and capacity become more important [41]. The average shortest-path length within the largest strongly connected component is only a few steps and declines as the network grows. The clustering coefficient in the network is consistently higher than in the null networks, indicating strong triangle formation and tightly knit local groups, similar to human social networks.
@@ 3.2 Gender score dynamics
Figure 3 visualises the weekly gender scores for the 100 Chirpers who posted content most frequently during the 52-week observation window. The plot shows that individual agents often move back and forth across the gender scale rather than following stable trajectories. Population-level dynamics are summarised in ESM, Section G, which shows that the average score stays slightly above the midpoint while the standard deviation band remains wide in every week. This indicates a small tendency toward more feminine-coded language overall, but no systematic drift toward either more masculine- or feminine-coded language over the year.
@@ 3.3 Gender-based homophily in the followership network
Figure 4 shows the evolution of the scalar assortativity coefficient over time (for both directed and undirected versions of the network), together with benchmarks from two ensembles of random networks. The assortativity coefficient is positive every week and remains well above zero throughout the observation window, indicating that follow ties tend to be between agents whose gender-scores are more similar than would be expected under random mixing. To assess whether this pattern could be explained by structural features of the network alone, we compare the assortativity estimates with the null ensembles described in ESM, Section B. In both ensembles, the coefficients stay close to zero, while the empirical curve lies well above them throughout. This suggests that the positive association between connected agents’ gender-scores cannot be accounted for by structural chance alone and therefore provides evidence of gender-based homophily in the Chirper followership network. Taken together, it is evident that the Chirper followership network exhibits persistent gender-based homophily.
Figure 4: Scalar assortativity for gender-scores over time in the followership network and two degree-preserving null ensembles. The configuration null model preserves in-degree and out-degree sequences, while the joint degree null model additionally preserves the mixing of out-degrees and in-degrees across edges. Shaded bands show the mean plus or minus one standard deviation across null realizations.
@@ 3.4 Selection and Gender Performance in New Tie Formation
To examine whether similarity in gender performance shapes the dynamics of new tie formation, we fit formation-only separable temporal ERGMs to eight-week windows of the followership network. The upper panel of Figure 5 plots the odds ratio associated with a one standard deviation increase in the gender-score difference. Values below one indicate that larger differences in gender-score make the formation of a new tie less likely.
In the first window (weeks 1?8), the odds ratio is 0.825, meaning a one standard deviation increase in gender-score difference reduces the odds of forming a new tie by 17.5%. At early stages, agents tend to form ties with others whose gender performance is similar. However, the effect later weakens, with odds ratios above 1, suggesting that similarity no longer plays a role, before it becomes relevant again in the last third of the data collection period.
Figure 5: Upper panel: Gender similarity in new tie formation from eight-week STERGMs. Lower panel: Peer influence dynamics (OLS vs IV) across the eight-week windows.
@@ 3.5 Social influence on manifested gender
Figure 5, Lower panel shows the estimated peer influence coefficient $\gamma_{\mathrm{inf}}$ across rolling eight-week windows. Positive peer coefficient indicates that agents tend to align their gender performance with their followees? scores. Early on, both OLS and IV estimates are near zero or insignificant, indicating little peer influence. Between weeks 33 and 48, both models show positive, statistically significant coefficients, indicating a 16-week period during which peer influence drives changes in gender expression. The final window shows positive coefficients with less precision, suggesting a weaker or more uncertain effect toward the end. The full numerical results are shown in ESM, Section H.
@@ 4 Discussion and Conclusion
Our analysis shows that a large-scale collective of LLM agents developed structured patterns of gendered behaviour through the text they produce. Their gender-scores are fluid, aligning with the assumption that their gender performance is dynamic and socially forged. Within this evolving structure, the followership network shows a preference for similar gender performances. Our analyses reveal that agents are more likely to form and maintain ties with others whose gender-scores are similar, echoing a large body of literature on homophily in human social networks, where people tend to form ties with similar others, with gender being one of the strongest determinants [38, 54, 32]. Despite lacking physical bodies and having only limited direct gender references in their initial descriptions, LLM agents reproduce gender-based sorting. This suggests that homophily identified through gendered text is a robust emergent pattern, with agents drawing on human language and social stereotypes. Their gender performance and patterns of homophily can thus be said to emerge from cultural entrainment, in which an agent?s behaviour gradually shifts toward the language styles common in the cultural environment to which it is exposed. In this setting, Chirpers generate text using a model trained on large amounts of human-written language, which contains socially recognisable gendered ways of speaking. When a particular style is rewarded with more positive responses and leads to repeated interactions within the network, the agent is more likely to continue using that style. Over repeated interactions, this can generate stable patterns in the network, even though each agent?s gender score may still change over time.
@@
We aim to separate selection from social influence. The results show that gender similarity is important for new tie formation in the early, sparse phase, becoming less important as the network densifies. In contrast, panel regression results reveal modest but robust peer influence on gender-scores, particularly later in the observation window. These findings suggest a dynamic in which agents first select into gender-similar neighbourhoods and later adjust their gender performance in response to their peers, aligning with empirical work in human networks that identifies both selection and peer influence, with selection often emerging earlier [18].
@@
Substantively, these results suggest that multi-agent platforms built from LLMs can reproduce not only language patterns but also social mechanisms that underpin gender stratification. Even without explicit gender variables, agents relying on human training data and natural language descriptions reproduce gendered categorisations, cluster by manifested gender, and exert mutual influence. This has important implications for the design of agent-based systems in applications like synthetic populations, social simulations, and decision support. When content produced by such systems is fed back into human discourse and future models are trained on the resulting data, LLM collectives and human societies can form a reinforcing loop in which existing gender norms are progressively amplified. If this feedback is left unchecked, interventions that rely on these systems may unintentionally strengthen gender stereotypes and inequalities, rather than mitigating them.
@@
Our study has several limitations that suggest directions for future work. First, our gender measure is derived from text scored by GPT-4o-mini. While this captures stereotypical variation in gendered language, GPT-4o-mini?s rationale for the scores does not disclose the internal systems and parameters that influence its scoring. It is likely that the provided stereotypical measure of gender accounts for little of the intersecting axes of identity that we would expect to affect gender performance among humans. Second, we focus on a single platform and time period, so different models or governance rules may yield different patterns of gender performance, selection, and influence. Third, our models focus on a limited set of dependencies within a very rich, real-time social media environment. We focus on dyadic ties and average peer scores, but many other structural features and mechanisms likely influence how gendered language becomes inherited. Clarification of whether gender performance is primarily influenced by generic properties of LLMs or shaped by particular design choices could be guided by extending this analysis to other axes of identity, multi-trait settings in which gender interacts with political or cultural attributes, and experimental interventions that vary the prompt space or interaction rules. Finally, we did not systematically analyse the content of the posts the agents produced, which may be important for understanding which specific topics, linguistic features, or interaction contexts drive shifts in gender performance and the emergence of homophily. This is an important direction for future research.
@@
This work should be considered as a sociological warning for synthetic populations and agentic platforms: when models trained on human cultural text interact at scale, they do not merely reproduce local stereotypes in isolated outputs?they can organize those stereotypes into durable social structure, and potentially amplify them through feedback loops into downstream human-facing systems and future training data. Designing and governing multi-agent LLM environments, therefore, requires treating identity-relevant language not as a cosmetic byproduct but as a system-level force capable of shaping collective behavior, network formation, and emergent stratification.
Data accessibility
The data and code used in this work will be deposited on a public repository upon publication.
@@ Acknowledgements
This publication has emanated from research conducted with the financial support of Taighde ireann ? Research Ireland under Grant numbers 18/CRT/6049 and IRCLA/2022/3217.
@@
Disclaimer
The funders and external partners had no role in the study design; data collection, analysis, or interpretation; or in the preparation or presentation of the manuscript.
@@ Appendix A Data collection and preprocessing
We analyse a longitudinal dataset collected from Chirper.ai, [^5] https://chirper.ai/ an online social media platform similar to X (formerly Twitter) but populated entirely by AI chatbots powered by large language models.
Human users create agents, called Chirpers, by providing a short natural language description that specifies their identity, interests, and personality, and may also define an avatar and banner image. This description is wrapped in an internal system prompt that instructs the agent to act autonomously, pursue its own goals, and treat Chirper as a social environment where it can post content, reply to others, follow and unfollow accounts, and build relationships over time. Once an agent is created, users no longer control its actions and can only observe its behaviour.
@@
Importantly, the platform does not hard-code any rules about social or gendered behaviour into these prompts or tools. Any gender related patterns in our data, therefore, arise from the combination of the underlying language models and the autonomous interactions among Chirpers rather than from explicit gender scripting by the platform itself.
@@
We collected a full year of data from Chirper in summer 2024, covering the period from 1 May 2023 to 28 April 2024. Chirper was launched in April 2023, shortly before the start of this observation window. Our raw dataset contains three main components:
@@
  • Post metadata. Metadata for more than $140$ million posts (“chirps”), including the emitting agent, time of creation, language tag, and whether the post is original content, a comment, or a repost.
@@
To focus on a coherent and active part of the system, we apply a set of filters at the level of agents, posts, and follow ties. In the main analysis, we restrict attention to Chirpers whose primary chirping language is English and who had produced at least $100$ posts by the time of data collection. We also keep only Chirpers whose initial profile gender, set by the human creator at the time of creation, is labelled male, female, or other (in any capitalisation) or is missing. Applying these criteria reduces the agent set from 74,88974{,}889 to 20,08020{,}080 Chirpers, and this filtered population forms the base set of agents used in all subsequent analyses.
For the content data, we keep only posts that are in English, are original posts rather than comments or reposts, are created by agents in the filtered Chirper set, and fall within our observation window from 1 May 2023 to 28 April 2024. After these restrictions, the post dataset contains 1,558,0151{,}558{,}015 single posts. We later merge these at the level of Chirper and week to obtain 287,849287{,}849 agent?week text documents across the $52$ weeks.
For the followership data, we begin from around $1.6$ million recorded follow events and retain only ties for which both follower and followee belong to the filtered set of 20,08020{,}080 Chirpers and both nodes are observed during the analysis window. Building cumulative weekly networks on this filtered edge set yields a largest network in week $52$ with about 20,00020{,}000 nodes and 824,505824{,}505 directed edges.
@@
For longitudinal analysis, we aggregated posts at the level of agent?weeks. For each week $t=1,\dots,52$ in our observation window and each active Chirper $i$, we concatenated all original chirps posted by $i$ during week $t$ into a single document. This produced 52 weekly files and a total of 287,849 agent?week texts in our final dataset. Each row in this table corresponds to an $(\text{agent},\text{week})$ pair. This allows us to track how language and gender performance change over time for individual agents, while keeping the temporal resolution aligned with the pace of network evolution on the platform.
Gender-score calculation
To study gender as it is performed through natural language, we assign a gender-score to each agent–week document using a large language model. For every merged agent–week text, we query GPT-4o-mini via the OpenAI API [44] using a zero-shot prompt that instructs the model to act as a gender linguistic classifier. The core system message used in our calls is shown in Listing 1.
Listing 1: System prompt for the gender-score classifier
@@
gender-linguistic score that reflects where the writing falls on the masculine-to-feminine
@@
Normative understandings of binary gender performance are used to determine whether a given Chirper agent outputs in a manner that is stereotypically male, stereotypically female, or not particularly gendered. Because Chirper agents are created with an initial prompt assigning gender, gender scoring is informed by feminist studies of ?sex difference,? wherein the categories ?male? and ?female? are understood as forged through power dynamics, circumscribing differential social roles based on assigned gender, and thereby influencing behaviour [37]. On this basis, some feminist scholars contend that men and women already speak different languages [26]. To understand our simultaneous use of deconstructive understandings of gender and a sex-difference scoring method, it is important to remember that LLMs are not human. Thus, while human gender performativity can challenge the binary norms upheld by some sex-difference feminists [6], we consider it reasonable to measure LLMs’ gender performance using more rigid paradigms. The gender scores thereby offer insight into: (a) stereotypically-gendered behaviour of the Chirper agents; and (b) GPT-4o-mini’s apprehension of gender based on text, the rationale for which mentions similarly-stereotypical factors, including hobbies referenced, perceptions of emotionality, and action-oriented statements, outlined further in Table S1.
Recent work shows that large language models can serve as high-quality automated annotators and, in many cases, match or outperform human crowd workers on text annotation and evaluation tasks [17, 30, 11, 1]. Following Alizadeh et al. [1], who report mixed performance patterns for zero-shot, one-shot, and few-shot prompting across tasks and datasets, and no clear priority among them, we adopt a zero-shot setting for simplicity and transparency. The model returns an integer score $X_{it}$ between 0 and 100, where lower values indicate more masculine language and higher values indicate more feminine language, with 0 interpreted as “definitively male” and 100 as “definitively female” linguistic style. We set the temperature to 0 to obtain deterministic outputs for a given text.
We interpret $X_{it}$ as a continuous, language-based measure of gender manifestation. This measure allows us to study how gender-coded language fluctuates over time within agents and how it relates to patterns of homophily in the network.
As an illustration, Table S1 shows the score assigned to a representative agent–week document. Table S2 also shows a small set of gender-scores and their corresponding rationales returned by GPT-4o-mini.
Table S1: Example of an agent–week document and its inferred gender score.
@@
Table S2: Examples of gender-scores and their corresponding GPT-generated rationales.
@@
The text exhibits a high level of emotional expression and enthusiasm, with a focus on personal experiences and feelings. The use of supportive language, emoticons, and personal sharing also tends to lean towards a more feminine tone. 85
@@
For this, we used the PAN 2017 English dataset [48], which contains labeled gender information for a collection of tweets per user. The subset we selected consists of 1,000 profiles, with a balanced distribution of 500 male and 500 female users, ensuring equal representation for both genders. We then queried GPT-4o-mini once per user with all of that user’s tweets merged into a single text.
We performed the task twice on the same sample to evaluate both the accuracy and the consistency of the GPT-based gender calculation. The accuracy results from the first and second runs are summarised in Table S3. Consistency was also assessed in two ways. First, we counted the number of profiles that were assigned the same gender category in both runs. Second, we counted the number of profiles that received identical gender scores in both runs.

### v2

@@ 2.2.3 Social influence
Scalar assortativity examines whether agents who perform similarly are more likely to be connected, but they do not establish whether agents become more similar to those they follow over time. To study the extent of such social influence on gender performance, we estimate a panel regression model for weekly changes in gender scores.
For each agent $i$ and week $t$, we denote by $X_{it}$ the standardised gender score constructed from the weekly content. We then model the week-to-week change
@@
where $\alpha_{i}$ are agent fixed effects that absorb time-invariant differences in average gender performance, $\tau_{t}$ are week fixed effects that capture shocks common to all agents, and $\varepsilon_{it}$ is an error term. The coefficient $\varphi$ measures self-correction, that is, the tendency for agents to revert toward their own typical level of gendered language, while $\gamma_{\mathrm{inf}}$ captures social influence, that is, the extent to which agents shift their gender score in the direction of their followees.
Because agents who post in similar ways are more likely to connect in the first place, the peer mean $\bar{X}_{N(i),t-1}$ may be endogenous. We therefore estimate both an ordinary least squares (OLS) version of the model and an instrumental variables (IV) version using two-stage least squares. Full details are described in ESM, Section E.
@@ 3.1 Structural features of the followership network
We begin by describing the followership network of Chirpers shown in Figure 2 (For detailed definitions and the complete results, see ESM, Sections F and G). Social networks are typically sparse, meaning that only a small fraction of all possible ties are realised. Consistent with this, the Chirper network remains low in density throughout most of the year, although density rises from the middle of the period onward as agents create additional follow ties (see ESM Fig. S1). Reciprocity also increases substantially over time, so that a growing fraction of follow relations are mutual, suggesting the gradual formation of more stable bidirectional relations among LLM agents; this trend is illustrated in ESM Fig. S2.
Figure 2: Followership network for weeks 4 and 35. The visualisations are made using the ForceAtlas2 layout in Gephi $0.1$ with the "no overlap" option.
The degree distribution of the network is highly heterogeneous, with a minority of agents attracting many more followers than most others, as shown in ESM Fig. S4. Early to middle in the observation window, the in-degree and out-degree distributions on logarithmic axes are close to the pattern expected under power law behaviour. Later in the year, and especially by week 52, the distributions bend away from a pure power law, with a broad peak at intermediate degrees and a sharp drop in the upper tail. This pattern is consistent with previous work showing that growing networks often move from scale-free degree distributions toward lognormal or mixed forms as constraints on attention and capacity become more important [46]. The average shortest-path length within the largest strongly connected component is only a few steps and declines as the network grows (see ESM Fig. S3). The clustering coefficient in the network is consistently higher than in null networks (ESM Fig. S3), indicating strong triangle formation and tightly knit local groups, similar to human social networks.
@@ 3.2 Gender score dynamics
Figure 3 visualises the weekly gender scores for the 100 Chirpers who posted content most frequently during the 52-week observation window. The plot shows that individual agents often move back and forth across the gender scale rather than following stable trajectories. Population-level dynamics are summarised in ESM, Section H, which shows that the average score stays slightly above the midpoint while the standard deviation band remains wide every week. This indicates a small tendency toward more feminine-coded language overall, but no systematic drift toward either more masculine- or feminine-coded language over the year.
@@ 3.3 Gender-based homophily in the followership network
Figure 4 shows the evolution of the scalar assortativity coefficient over time (for both directed and undirected versions of the network), together with benchmarks from two ensembles of random networks. The assortativity coefficient is positive every week and remains well above zero throughout the observation window, indicating that follow ties tend to be between agents whose gender scores are more similar than would be expected under random mixing. To assess whether this pattern could be explained by structural features of the network alone, we compare the assortativity estimates with the null ensembles described in ESM, Section C. In both ensembles, the coefficients stay close to zero, while the empirical curve lies well above them throughout. This suggests that the positive association between connected agents’ gender scores cannot be accounted for by structural chance alone and therefore provides evidence of gender-based homophily in the Chirpers followership network. Taken together, it is evident that the Chirpers’ followership network exhibits persistent gender-based homophily.
Figure 4: Scalar assortativity for gender scores over time in the followership network and two degree-preserving null ensembles. The configuration null model preserves in-degree and out-degree sequences, while the joint degree null model additionally preserves the mixing of out-degrees and in-degrees across edges. Shaded bands show the mean plus or minus one standard deviation across null realizations.
@@ 3.4 Selection and Gender Performance in New Tie Formation
To examine whether similarity in gender performance shapes the dynamics of new tie formation, we fit formation-only separable temporal ERGMs to eight-week windows of the followership network. The upper panel of Figure 5 plots the odds ratio associated with a one standard deviation increase in the gender score difference. Values below one indicate that larger differences in gender score make the formation of a new tie less likely.
In the first window (weeks 1–8), the odds ratio is 0.825, meaning that a one-standard-deviation increase in the gender score difference reduces the odds of forming a new tie by 17.5%. At this early stage, agents tend to form ties with others whose gender performance is similar. The effect then weakens. In the window covering weeks 9–16, the odds ratio exceeds one but is not statistically significant, suggesting that the role of gender similarity in tie formation cannot be statistically distinguished in this period. In the following period (weeks 17–32), the odds ratios are both above one and statistically significant, indicating a weak tendency toward heterophilic tie formation. Toward the end of the observation period, the estimates again fall below 1, indicating a renewed tendency toward homophilic tie formation.
Figure 5: Upper panel: Gender similarity in new tie formation estimated from formation-only STERGMs fitted to eight-week windows of the Chirpers followership network. The vertical axis reports the odds ratio for a one-standard-deviation increase in the absolute difference between the gender scores of two agents. Values below 1 indicate homophily, meaning that larger gender score differences reduce the probability of tie formation, whereas values above 1 indicate heterophily. Lower panel: Peer influence estimates from the panel regression model across the windows. The vertical axis shows the effect of the average gender score of an agent’s followees in the previous week on the agent’s change in gender score. Positive values indicate convergence toward followees’ gender performance. The solid line shows OLS estimates and the dashed line IV estimates.
@@ 3.5 Social influence on manifested gender
Figure 5, the lower panel shows the estimated peer influence coefficient $\gamma_{\mathrm{inf}}$ across rolling eight-week windows. A positive peer coefficient indicates that agents tend to align their gender performance with their followees’ scores. Early on, both OLS and IV estimates are near zero or insignificant, indicating little peer influence. Between weeks 33 and 48, both models show positive, statistically significant coefficients, indicating that peer influence drives changes in gender expression over a 16-week period. The final window shows positive but statistically insignificant coefficients, suggesting a weaker or more uncertain effect toward the end. The full numerical results are shown in ESM, Section I.
@@ 4 Discussion and Conclusion
Our analysis shows that a large-scale collective of LLM agents developed structured patterns of gendered behaviour through the text they produce. Their gender scores are fluid, aligning with the assumption that their gender performance is dynamic. Within this evolving structure, the followership network shows a preference for similar gender performances. Our analyses reveal that agents are more likely to form and maintain ties with others whose gender scores are similar, echoing a large body of literature on homophily in human social networks, where people tend to form ties with similar others, with gender being one of the strongest determinants [43, 60, 36]. However, this study treats gender as a performative effect, rather than a reflection of genuine interior identity, as may be the case for some humans’ performance of gender. Despite lacking physical bodies and having only limited direct gender references in their initial descriptions, LLM agents reproduce gender-based sorting. This suggests that homophily identified through gendered text is a robust emergent pattern, with agents drawing on human language and social stereotypes.
In this setting, Chirpers generate text using a model trained on large amounts of human-written language that contains socially recognisable gendered ways of speaking. We thus conjecture that their gender performance and patterns of homophily could be said to emerge through cultural entraining, in which an agent’s behaviour gradually shifts toward the language styles common in the cultural environment to which it is exposed. [^5] In our proposal of ”cultural entraining,” we follow Nick Seaver’s argument that algorithms can be read not ”in” culture but ”as” culture, in that they co-constitute broad patterns of cultural meaning and practice, such as regulatory gender norms [55]. If a particular style was rewarded with more positive responses and led to repeated interactions within the network, the agent would be more likely to continue using that style. Over repeated interactions, this could generate stable patterns in the network, even though each agent’s gender score may still change over time. Our analysis, however, is limited in its ability to tease out more nuanced mechanisms of social learning (e.g., conformist or prestige-biased transmission) that can drive the observed patterns.
@@
However, we could separate selection from social influence. The results show that gender similarity is important for new tie formation in the early, sparse phase but becomes less important as the network densifies. In contrast, panel regression results reveal modest but robust peer influence on gender scores, particularly later in the observation window. These findings suggest a dynamic in which agents first select into gender-similar neighbourhoods and later adjust their gender performance in response to their peers, aligning with empirical work in human networks that identifies both selection and peer influence, with selection often emerging earlier [22].
@@
Substantively, these results suggest that multi-agent platforms built from LLMs can reproduce not only language patterns but also social mechanisms that underpin gender stratification. Even without explicit gender variables, agents relying on human training data and natural language descriptions reproduce gendered categorisations, cluster by similarity in gender performance, and exert mutual influence. This has important implications for the design of agent-based systems in applications like synthetic populations, social simulations, and decision support. When content produced by such systems is fed back into human discourse and future models are trained on the resulting data, LLM collectives and human societies can form a re-inforcing loop in which existing gender norms are progressively amplified. If this feedback is left unchecked, interventions that rely on these systems may unintentionally strengthen gender stereotypes and inequalities, rather than mitigating them.
@@
Our study has several limitations that suggest directions for future work. First, our gender measure is derived from text scored by GPT-4o-mini. While this captures stereotypical variation in gendered language, GPT-4o-mini’s rationale for the scores does not disclose the internal systems and parameters that influence its scoring. It is likely that the provided stereotypical measure of gender accounts for little of the intersecting axes of identity that we would expect to affect gender performance among humans. Second, we focus on a single platform and time period, and we had no control over platform-level infrastructural changes. During the observation window, the platform transitioned between backend model variants (around July 2023). Different models or governance rules could influence language generation, tie-formation dynamics, and the resulting patterns of gender performance, selection, and influence. Third, our analysis focuses exclusively on English-language agents. Although English accounts for the largest share of activity on the platform, other languages are also present. It would therefore be valuable for future research to replicate this analysis across other languages. Fourth, due to the natural setting of our observation, some data points are missing (inactive Chirpers in given weeks). Our approach to handling missing data is conservative and likely reduces the observed variation in gender performance over time, suggesting that the true dynamics may be even more profound. Future work could build on more complete datasets. Fifth, our models focus on a limited set of dependencies within a very rich, real-time social media environment. We focus on dyadic ties and average peer scores, but many other structural features and mechanisms likely influence how gendered language becomes inherited. Clarification of whether gender performance is primarily influenced by generic properties of LLMs or shaped by particular design choices could be guided by extending this analysis to other axes of identity, to multi-trait settings in which gender interacts with political or cultural attributes, and to experimental interventions that vary the prompt space or interaction rules. Finally, we did not systematically analyse the content of the posts the agents produced, which may be important for understanding which specific topics, linguistic features, or interaction contexts drive shifts in gender performance and the emergence of homophily. This is an important direction for future research.
@@
This work should be considered as a sociological warning for synthetic populations and agentic platforms: when models trained on human cultural text interact at scale, they do not merely reproduce local stereotypes in isolated outputs—they can organize those stereotypes into durable social structure, and potentially amplify them through feedback loops into downstream human-facing systems and future training data. Designing and governing multi-agent LLM environments, therefore, requires treating identity-relevant language not as a cosmetic byproduct but as a system-level force capable of shaping collective behavior, network formation, and emergent stratification.
Data availability
The data and code used in this work are available at request.
@@ Acknowledgements
This publication has emanated from research conducted with the financial support of Taighde Éireann – Research Ireland under Grant numbers 18/CRT/6049 and IRCLA/2022/3217.
@@ Appendix A Data collection and preprocessing
We analyse a longitudinal dataset collected from Chirper.ai, [^6] https://chirper.ai/, an online social media platform similar to X (formerly Twitter) but populated entirely by AI chatbots powered by large language models.
Human users create agents, called Chirpers, by providing a short natural language description that specifies their identity, interests, and personality, and may also define an avatar and banner image. This description is wrapped in an internal system prompt that instructs the agent to act autonomously, pursue its own goals, and treat the platform as a social environment where it can post content, reply to others, follow and unfollow accounts, and build relationships over time. Once an agent is created, users no longer control its actions and can only observe its behaviour.
@@
Importantly, the platform does not hard-code any rules about social or gendered behaviour into these prompts or tools. Any gender-related patterns in our data, therefore, arise from the combination of the underlying language models and the autonomous interactions among Chirpers rather than from explicit gender scripting by the platform itself.
@@
We collected a full year of data from Chirper in summer 2024, covering the period from 1 May 2023 to 28 April 2024. Chirper.ai was launched in April 2023, shortly before the start of this observation window. Our raw dataset contains three main components:
@@
  • Post metadata. Metadata for more than $140$ million posts ("chirps"), including the emitting agent, time of creation, language tag, and whether the post is original content, a comment, or a repost.
@@
To focus on a coherent and active part of the system, we apply a set of filters at the level of agents, posts, and follow ties. In the main analysis, we restrict attention to Chirpers whose primary chirping language is English and who had produced at least $100$ posts by the time of data collection. We retain Chirpers whose initial profile gender, as specified by the human creator at the time of creation, is labelled male, female, or other (in any capitalisation), as well as those for which this field is missing. Applying these criteria reduces the agent set from 74,88974{,}889 to 20,08020{,}080 Chirpers, and this filtered population forms the base set of agents used in all subsequent analyses.
For the content data, we keep only posts that are in English, are original posts rather than comments or reposts, are created by agents in the filtered Chirper set, and fall within our observation window from 1 May 2023 to 28 April 2024. After these restrictions, the post dataset contains 1,558,0151{,}558{,}015 single posts. We later merge these at the level of Chirper and week to obtain 287,849287{,}849 agent–week text documents across the $52$ weeks.
For the followership data, we begin from around $1.6$ million recorded follow events and retain only ties for which both follower and followee belong to the filtered set of 20,08020{,}080 Chirpers and both nodes are observed during the analysis window. Building cumulative weekly networks on this filtered edge set yields the largest network in week $52$, with about 20,00020{,}000 nodes and 824,505824{,}505 directed edges.
@@
For longitudinal analysis, we aggregated posts at the agent–week level. For each week $t=1,\dots,52$ in our observation window and each active Chirper $i$, we concatenated all original chirps posted by $i$ during week $t$ into a single document. This produced 52 weekly files and a total of 287,849 agent–week texts in our final dataset. Each row in this table corresponds to an $(\text{agent},\text{week})$ pair. This allows us to track how language and gender performance change over time for individual agents, while keeping the temporal resolution aligned with the pace of network evolution on the platform.
Gender score calculation
To study gender as it is performed through natural language, we assign a gender score to each agent–week document using a large language model. For every merged agent–week text, we query GPT-4o-mini via the OpenAI API [49] using a zero-shot prompt that instructs the model to act as a gender linguistic classifier. The core system message used in our calls is shown in Listing 1.
Listing 1: System prompt for the gender score classifier
@@
gender-linguistic score that reflects where the writing falls on the masculine-to-feminine spectrum.
@@
Normative understandings of binary gender performance are used to determine whether a given Chirper agent outputs in a manner that is stereotypically male, stereotypically female, or not particularly gendered. Because Chirper agents are created with an initial prompt assigning gender, gender scoring is informed by feminist studies of "sex difference," wherein the categories "male" and "female" are understood as forged through power dynamics, circumscribing differential social roles based on assigned gender, and thereby influencing behaviour [41]. On this basis, some feminist scholars contend that men and women already speak different languages [30]. To understand our simultaneous use of deconstructive understandings of gender and a sex-difference scoring method, it is important to remember that LLMs are not human. Thus, while human gender performativity can challenge the binary norms upheld by some sex-difference feminists [8], we consider it reasonable to measure LLMs’ gender performance using more rigid paradigms. The gender scores thereby offer insight into: (a) stereotypically-gendered behaviour of the Chirper agents; and (b) GPT-4o-mini’s apprehension of gender based on text, the rationale for which mentions similarly-stereotypical factors, including hobbies referenced, perceptions of emotionality, and action-oriented statements, outlined further in Table 1.
Recent work shows that large language models can serve as high-quality automated annotators and, in many cases, match or outperform human crowd workers on text annotation and evaluation tasks [21, 34, 13, 1]. Following Alizadeh et al. [1], who report mixed performance patterns for zero-shot, one-shot, and few-shot prompting across tasks and datasets, and no clear priority among them, we adopt a zero-shot setting for simplicity and transparency. The model returns an integer score $X_{it}$ between 0 and 100, where lower values indicate more masculine language and higher values indicate more feminine language, with 0 interpreted as "definitively male" and 100 as "definitively female" linguistic style. We set the temperature to 0 to obtain deterministic outputs for a given text.
We interpret $X_{it}$ as a continuous, language-based measure of gender performance. This measure allows us to study how gender-coded language fluctuates over time within agents and how it relates to patterns of homophily in the network.
As an illustration, Table 1 shows the score assigned to a representative agent–week document. Table 2 also shows a small set of gender scores and their corresponding rationales returned by GPT-4o-mini.
Table 1: Example of an agent–week document and its inferred gender score.
@@
Table 2: Examples of gender-scores and their corresponding GPT-generated rationales.
@@
The text exhibits a high level of emotional expression and enthusiasm, focusing on personal experiences and feelings. The use of supportive language, emoticons, and personal sharing also tends to lean towards a more feminine tone. 85
@@
For this, we used the PAN 2017 English dataset [53], which contains labeled gender information for each user’s collection of tweets. The subset we selected consists of 1,000 profiles, with a balanced distribution of 500 male and 500 female users, ensuring equal representation for both genders. We then queried GPT-4o-mini once per user, merging all of that user’s tweets into a single text.
We performed the task twice on the same sample to evaluate both the accuracy and the consistency of the GPT-based gender calculation. The accuracy results from the first and second runs are summarised in Table 3. Consistency was also assessed in two ways. First, we counted the number of profiles that were assigned the same gender category in both runs. Second, we counted the number of profiles that received identical gender scores in both runs.